# Can the 3 neutrino masses really be found using SN 1987A data?

### v1

@@ IV Fitting 30 data points to 3 lines
Strikingly, when the SN 1987A data is plotted (see Fig.1) every one of the 30 neutrinos lies on one of three straight lines through the origin, and very few are next to two lines. The least square fit of the 30 data points to three lines through the origin has two free parameters. The first parameter is $S$ which adjusts the trial slopes of the three straight lines in Fig. 1 simultaneously and the second free parameter H is the size of the constant horizontal error bar on each data point in Fig.1. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square. However, only a few of the data points near the origin are within $1\sigma$ of more than one straight line. The good fit shown in Fig 1 has a chi square of 25.4. Given 28 degrees of freedom we find $p=61\%.$ The excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=18\%$.) For a fuller desccription of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig. 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot. Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is no more than 1-2 seconds.
@@
Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to three straight lines that pass through the origin. Error bars in both horizontal and vertical are included in the fit. The best fit has $\chi^{2}=25.0$ for 28 dof, which yields $p=61\%$. The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq. 5. Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity.
@@
The three masses corresponding to the slopes of the three straight lines are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.7eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line. The big difference from the author’s previously published values of this fit is in the value of $m_{2}$ because of the past omission of horizontal error bars, which are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line. We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.7eV/c^{2}$ and the $2.7eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration, as well as its non-observation by KATRIN. (Serebrov2021, ; KATRIN2025, ). Of course, our $m_{1},m_{2},m_{3},$ however, are not sterile neutrino masses, but the three active neutrino masses.
@@
Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points. Obviously for $H>>5$ sec error bars the fit has $p\approx 100\%$
V Discussion
@@
In this section we discuss eleven issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply three specifc neutrino masses, and it is not just a mirage or background fluctuation.
V.1 Why not not exclude the LSD data?
@@
Initially, the author rejected the LSD data. This rejection was not because it proviided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this and all the other reasons many people have for rejecting the LSD data can be answered, there really is no valid reason not to keep it.
V.2 Why are the LSD data bunched in time?
@@
If the LSD burst of 5 neutrinos is genuine, it looks strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data has a kilosecond time scale. If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a bunch of events in a seconds-long time interval.
V.3 Why do the LSD neutrinos have the same E?
@@
The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines. Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it. That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and further evidence related to the observed spectrum of MeV $\gamma$-rays from the galactic center. (Ehrlich2021, )
V.4 Alternate explanations for the LSD neutrinos?
@@
All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
@@
  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section 6.
@@
  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still creates the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
V.5 The meaning of the horizontal error bars?
@@
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arriival time randomly shifted by a time up to that amount, and a horizonal error bar of that amount or greater would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 1-2 sec.
V.6 Does a $\Delta t=1-2$ sec agree with simulations?
@@
We can find the actual size of the time window using a simulation of neutrino emission. In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 3. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background. These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 1-2 sec time window, and that our use of 2 sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
V.7 Why don’t late-emitted $\nu$’s mess up the fit?
First, because based on Fig.3 there are not very many of these, perhaps no more than 15-20% of the total for $t>2$ sec, and second because of the late neutrinos associated with the shallower sloped line in Fig. 1, a shift of the points horizontally even by quite a few seconds on a shallow sloped line has little effect on the goodness of its fit to the line, which depends mainly on the vertical error bar. Thus, for example, a data point lying one sigma away vertically from that line could still be less than one sigma away if the point was shifted as much as five seconds horizontally.
@@ V.7 Why don’t late-emitted $\nu$’s mess up the fit?
Figure 3: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the different time scales used for $t<1$ sec and $t>1$ sec. The fraction of the integrated luminosity for $t<1$ is less than it seems because of this scale change. Figure is used with permission of publisher
V.8 Two really big vertical error bars?
Two data points occurring at the largest t in Fig. 1 appear anomalous particularly because of their very large error bars compared to all others. This is the result of error propagation and the fact that on a plot of $1/E^{2}$ versus t the size of the vertical error bar varies as the inverse cube of the energy E according to $\Delta(1/E^{2})=2\Delta E/E^{3}.$ In any case these two points degrade the fit very little because of their large error bars.
V.9 What if all three $m_{k}$ are all very small?

### v2

@@ III Fitting 30 data points to 3 lines
Thirty neutrinos is not a very large number, but the distinctiveness of a pattern counts for far more than the sheer number of events. Thus, the $\Omega^{-}$ particle which led to quarks was confirmed in 1964 at Brookhaven National Laboratory by observing just one single“Gold Standard” event. (Barnes1964, ) In the present case, the pattern found when the SN 1987A data is plotted (see Fig.1) is also very distinctive since every one of the 30 neutrinos from SN 1987A lies on or next to one of three straight lines through the origin as predicted, and only three points near the origin are unavoidably next to two lines. The least square fit of the 30 data points to three lines through the origin has three free parameters, $S_{j}$ the slopes of the three straight lines in Fig. 1 yielding 27 degrees of freedom. The parameter H, the size of the constant horizontal error bar on each data point in Fig.1 is also adjusted. Since each data point has both x and y error bars, we can then convert the x error into an ”effective” y error by multiplying the x uncertainty by the slope of the line, and combine it with the y uncertainty in quadrature. If a data point lies close to two straight lines, we simply choose the line that gives the smaller contribution to chi square. Only a few of the data points lie beyond $1\sigma$ of a straight line.
The good fit shown in Fig 1 has a chi square of 25.4. Given 27 degrees of freedom we find $p=55\%.$ This excellent best fit was for a horizontal error bar of two seconds on all data points, and the fit would still be acceptable but significantly worse for a one second horizontal error bar ($p=15\%$.) For a fuller description of how the $\chi^{2}$ fit probability depends on the size of the horizontal error bar see Fig. 2 We have seen that exactly simultaneous neutrino emissions requires the data must lie on one of three lines through the origin in a $1/E^{2}$ versus t plot. Therefore the goodness of fits using 1 and 2 sec horizontal error bars shows they will continue to lie on or near these lines if the time window of neutrino emissions is also no more than 1-2 seconds.
@@
Figure 1: Best fit of 30 data points $(t,1/E^{2}),$ one for each neutrino to one of three straight lines that pass through the origin. Error bars in both horizontal and vertical are included in the fit. The fit has $\chi^{2}=25.4$ for 27 dof, which yields $p=55\%$. The three best fit masses $m_{j}$ are obtained from the three best fit slopes according to Eq. 5. Note the different scales used for $t>0$ and $t<0$ The five LSD data points at t=17ksect=-17ksec have been slightly separated from each other for clarity. The meaning of the shaded central area is discussed in the text
@@
The three masses corresponding to the slopes of the three straight lines in Fig. 1 are $m_{1}=22.4eV/c^{2}$ and $m_{2}=2.70eV/c^{2}$ for the positive sloped lines and $m_{3}^{2}=-200,000eV^{2}/c^{4}$ for the $m^{2}<0$ line. The author first discovered this three line fit in 2013, when he interpreted the result in terms of a 3 + 3 model (three pairs of active stable neutrinos – see Fig.3.) The big difference from the author’s previously published values of this fit (Ehrlich2013, ; Ehrlich2019, ) is in the value of $m_{2}$ because of the past omission of horizontal error bars. Those error bars are crucial for fitting the nearly vertical $m_{2}$ line and much less so for the $m_{1}$ line. We are not unaware of the interesting numerical coincidence of the value of $m_{2}=2.70eV/c^{2}$ and the $2.70eV/c^{2}$ sterile neutrino mass reported by the Neutrino-4 collaboration (both equal to within $1\%$), as well as its non-observation by KATRIN, (Serebrov2024, ; KATRIN2025, ) a matter discussed at length in section IV.9.
III.1 Fixing a limitation of sterile neutrino searches
Of course, the existence of sterile neutrinos is highly controversial, there are probably more negative search results than positive ones. Thus, neither MicroBooNE, nor IceCube, nor KATRIN nor DayaBay report any evidence for them. (Micro2025, ; Aartsen2020, ; KATRIN2025, ; An2024, ). Nevertheless, an omission in these and other experiments is that they have not sought evidence for the presence of multiple sterile neutrinos simultaneously. In effect, if you are only searching for one sterile neutrino when two or three are present they can mask each others presence. One way to avoid this problem is to search for two or three in the same fit to the data, and another way is to find a way to spatially separate any sterile neutrinos based on their masses, or find a place where nature causes this separation for us.
III.2 More evidence for the two $m^{2}>0$ masses
It is believed that dark matter holds together both our galaxy as well as clusters of galaxies. One would expect that if both heavy and light $m^{2}>0$ sterile neutrinos were present throughout space the heavy ones would preferentially congregate in an individual galaxy and the lighter ones would preferentially congregate an a galaxy cluster. M. H. Chan and the author have carried out fits to the “observed” dark matter distributions in both the Milky Way galaxy and in four clusters of galaxies. (Chan2014, ) The goal was to see what mass sterile neutrino would match the observed profile in each case. This was done by deriving an equation of state for the sterile neutrinos dependent on their mass, and then finding the dark matter radial profile in both the Milky Way and four groups of galaxies. Finally, the mass was varied to find the value that best fit the observed profile in each case. The results of these calculations showed that the two $m^{2}>0$ values from the 3-mass SN 1987A fit would do nicely with the heavier mass holding clusters together as would be expected. It seems unlikely one could find another case where nature is going to be so kind as to separate spatially different mass dark matter particles. However, as we shall see in section IV.9, KATRIN has a way to search for all three neutrino masses including the tachyonic one simultaneously in a single fit.
@@ III.2 More evidence for the two $m^{2}>0$ masses
Figure 2: The probability from the least squares fit to three straight lines through the origin versus the size of the horizontal error bars, assumed to be the same for all 30 data points. Note that there is no “best” fit to three straight lines unless one first fixes the size of the horizontal error bars because for $H>>5$ sec error bars the fit probability $p\approx 100\%$
Figure 3: A possible $3+3$ neutrino mass squared diagram postulated by the author in 2013, showing updated $m_{1}$ and $m_{2}$ values. There are 3 active-sterile doublets (solid and dotted lines) split by three $\Delta m^{2}$ The three $\Delta m^{2}$ are for solar, atmospheric and short baseline (sbl) oscillations. Diagram is not to scale. The $m_{2}$ value differs significantly from the previously found value, i.e. $(m_{2}=4.0eV/c^{2})$ rather than $2.70eV/c^{2},$ because of the use of horizontal error bars now used in the three mass fit
IV Discussion
@@
In this section we discuss 10 issues that need to be addressed before we can have any confidence that the excellent fit in fig.1 to three straight lines (three masses), including one $m^{2}<0,$ really does imply we can deduce the three neutrino masses from the SN 1987A data, and it is not just a mirage or background fluctuation.
IV.1 Why not exclude the LSD data?
@@
Initially, the author mistakenly rejected the LSD data. This rejection was not because it provided evidence for a $m^{2}<0$ neutrino, but because the idea of monochromatic neutrinos from a supernova seemed preposterous. However since this issue and all the other reasons many people mistakenly have for rejecting the LSD data can be answered, there is no valid reason not to keep LSD data, especially because with three neutrinos the mandated number of straight lines about which clustering should occur is three.
IV.2 Why are the LSD data so bunched in time?
@@
If the LSD burst of 5 neutrinos is genuine, it may look strange in Fig. 1 that while the other data points lie all along their respective lines the five LSD neutrinos are all bunched together in time at t=484t=-484 min or -17 ksec. The three straight line result would be more believable perhaps if the five neutrinos yielded data points at five widely spaced points on the negatively sloped line. But recall that the $t<0$ data has a kilosecond time scale. If the five LSD data points were spread out along the negative sloped line and not bunched in time they never could have been distinguished from background because researchers were looking for a concentration of neutrino events in a seconds-long time interval, i.e., a burst.
IV.3 Why do the LSD neutrinos have the same E?
@@
The apparent near-constant energy of the LSD burst remains mysterious even to the physicists who accept its existence, but it is natural if the 5 neutrinos (and all others) must lie on one of three straight lines. Moreover, there is a theoretical basis for the monochromaticity of the LSD burst, as shown in several of the author’s publications that have provided independent evidence for it. That evidence was in the form of a proposed model for an 8 MeV neutrino line from a supernova, and support for the model from the observed spectrum of MeV $\gamma$-rays from the galactic center. (Ehrlich2021, )
IV.4 Alternate explanations for the LSD neutrinos?
@@
Many phenomena in nature on the fringes of the known have alternative explanations. All the alternate explanations of the LSD neutrino burst being $m^{2}<0$ tachyons have their own problems. Thus,
@@
  • If it was due to a double bang the pattern in Fig. 1 should look very different as discussed in section IV.7.
@@
  • If it was due to 8 MeV dark matter as suggested in ref. (Ehrlich2021, ) this would still create the same LSD burst, so it is not really an alternative, but rather an explanation of the LSD neutrinos falling on the $m^{2}<0$ line.
IV.5 Best size of the horizontal error bars?
@@
The main way the good fit to three straight lines could be a mirage is that the neutrinos were not emitted simultaneously as our derivation requires, meaning that the size of the emissions time window was too large for the method to work. If all neutrinos were emitted simultaneously (a zero width time window) the recorded time of each neutrino would accurately represent its emission time, and there would be no need for having any horizontal error bars. If the time window were an amount $\Delta t$ then each neutrino might have its arrival time t randomly shifted to the right by a time up to that amount, and a horizonal error bar $H\approx\Delta t$ would give a good fit to three straight lines. Thus, we can find the size of the emission time window by seeing for what horizontal error bar width H, the goodness of the fit probability significantly improves. Clearly from Fig. 2 it is about 1-2 seconds.
IV.6 The time window for neutrino emissions?
@@
We can estimate the actual size of the time window for neutrino emissions using a simulation. In 2025 Bozza et al. have calculated the neutrino luminosity from SN 1987A using a new model they recently developed. (Bozza2025, ) They claim the analysis is in many ways more refined and accurate than previous ones with respect to the temporal structure of the neutrino luminosity. The functional form of their computed luminosity has 9 parameters that are determined by a least squares fit to the SN 1987A data, and their result is shown in Fig. 4. Note that the maximum neutrino emission occurs at about $T=100ms=0.1sec$ and the bulk of the emissions occur for $t<1$ second. Their best fit parameter values are all in accordance with current supernova simulations, and they find surprisingly that $95\%$ of the analyzed events (even those at $t>>1$ sec) are signal rather than background. These results all buttress our case that the SN 1987A neutrinos were mostly emitted during a 1-2 sec time window, and that our use of 2 sec for the optimal horizontal uncertainty in our least squares fit to three straight lines was justified. A 1 to 2 sec time window is apparently small enough not to badly degrade the fit of 30 neutrino data points to three straight lines.
@@
Figure 4: Computed temporal electron antineutrino luminosity L in units of $10^{53}erg/s$ for SN 1987A neutrinos from Bozza et al. (Bozza2025, ). Note the time scale change at $t=1$ sec. The fraction of the integrated luminosity for $t<1$ is less than it seems because of the scale change. Figure is used under creative commons license https://creativecommons.org/licenses/by/4.0
IV.7 Suppose neutrinos all have $m<<1eV/c^{2}$?